OpenStudy (anonymous):

A linear function has a slope of 7, and when x is 3, the value for y is 22. use a graph or table to figure out the equation of the function. show your work.

OpenStudy (anonymous):

Let y = mx + c be the equation of the linear function. Since the function has a slope of 7, we get m = 7. So, y = 7x + c When x=3, y=22: Substitute these values in the above function to find c: 22 = 7(3) + c c = ?